OpenStudy (anonymous):

2x – 5y = -3 -x + 2y = 3 i got: (-9,-3). am i right?

OpenStudy (e.cociuba):

yes ur rite!!!

OpenStudy (anonymous):

okay thank you :)

OpenStudy (anonymous):

2x - 5y = -3 - x + 2y = 3--> -x = -2y + 3 --> x = 2y - 3 now sub 2y - 3 in for x in the first equation 2(2y - 3) - 5y = - 3 4y - 6 - 5y = - 3 -y - 6 = - 3 -y = - 3 + 6 -y = 3 y = - 3 now sub - 3 in for y in either of original equations... -x + 2y = 3 -x + 2(-3) = 3 -x - 6 = 3 -x = 6 + 3 -x = 9 x = -9 check... 2x - 5y = -3 2(-9) - 5(-3) = - 3 -18 + 15 = - 3 -3 = -3 (correct) you are correct anthony...the answer is (-9,-3)

OpenStudy (e.cociuba):

np:)